OpenStudy (anonymous):

Select the ordered pair from the choices below that satisfies following system of equations: 2x - 10y = -14 4y = -x - 5 Answer (-1/2, 1) (-20, 3) (3, 2) (8, -4)

OpenStudy (texaschic101):

sub in your solutions and see which one is correct (-1/2,1) 2x - 10y = -14 2(-1/2) - 10(1) = -14 -1 -10 = -14 (we do not need to check the other equation because this solution does not work. (-20)(3) 2x - 10y = -14 2(-20) - 10(3) = -14 -40 - 30 = -14 -70 = -14 (will not work) (3,2) 2x - 10y = -14 2(3) - 10(2) = -14 6 - 20 = -14 (this works for this equation, now we need to check it with the other equation (3,2) 4y = -x - 5 4(2) = -3 - 5 8 = -8 (not correct) (8,4) 2x - 10y = -14 2(8) - 10(4) = -14 16 - 40 = -14 -24 = -14 (nope) I don't know....I can't get any of them to work

OpenStudy (texaschic101):

@Loser66 ...what did I do wrong ?

OpenStudy (loser66):

hehe, neither I

OpenStudy (texaschic101):

the (3,2) works in one equation, but not the other

OpenStudy (loser66):

We try all our best. hand off, friend. I don't replace, I solve for x and y from the system. My answer is not one of them.

OpenStudy (texaschic101):

I did that first....just solving by elimination...it didn't match any of your solutions

OpenStudy (loser66):

his solutions, not my solutions, friend

OpenStudy (texaschic101):

I meant his....I understand this is not your question....

OpenStudy (texaschic101):

Aphex....is everything in the problem, including solutions, correctly written

OpenStudy (texaschic101):

well....I tried

OpenStudy (austinl):

Something is clearly wrong.

OpenStudy (texaschic101):

I agree

OpenStudy (texaschic101):

x + 4y = -5 -->(-2) 2x - 10y = -14 ---------------- -2x - 8y = 10 (result of multiplying by -2) 2x - 10y = -14 ---------------add -18y = -4 y = 4/18 = 2/9 x + 4y = -5 x + 4(2/9) = -5 x + 8/9 = -5 x = -5 - 8/9 x = -45/9 - 8/9 x = - 53/9 check... 2x - 10y = -14 2(-53/9) - 10(2/9) = -14 -106/9 - 20/9 = -14 -126/9 = -14 -14 = -14 I got x = -53/9 and y = 2/9